OpenStudy (anonymous):

Lonnie places $950 in a savings account that earns 5.75% simple interest. Find the total amount in the account after four years.

OpenStudy (anonymous):

si=ptr/100 si=950*4*5.75/100 si=218.5 total amount 218.5+950=1168.5